# CLIMB: CLustering-based Iterative Data Mixture Bootstrapping for Language Model Pre-training

### v1

@@ 4.1 Implementation Details
5 Experimental Results
5.1 Comparison with Data Mixture Baselines
5.2 Comparison with SOTA LMs
Using the optimal data mixture identified by our method, we further investigate the effect of scaling up. Specifically, we used the same data mixture to train on 400B tokens and then compared the resulting model against state-of-the-art baselines. As shown in Table 2, CLIMB achieves the best performance among all sub-500M and sub-1.2B models. For example, when comparing models of similar scales (around 1B parameters), CLIMB consistently outperforms other baselines — including Llama-3.2 and AMD-OLMo — across the majority of the general reasoning benchmarks. In particular, it achieves the highest overall average score, surpassing the next-best method (i.e., Llama-3.2) by a noticeable margin (2.0%). Moreover, we introduced additional benchmarks (e.g., mmlu, gpqa, obqa, boolq, and race), and our model is consistently better than baseline models, which demonstrates that our method exhibits excellent generalization performance.
6 Analysis
Effects of Search Compute Budget. In the main experiments, we fix our total search budget (total compute) at 100%. Concretely, we perform three iterations of search with 64, 32, and 16 candidates evaluated in iterations 1, 2, and 3, respectively, giving a total of 112 searches. To understand how scaling search computes helps, we compare runs with greater total numbers of searches (e.g., 168, 224). Increasing the total number of searches allows the search procedure to more thoroughly explore possible data-mixture candidates each iteration. As shown in Table 3 (rows under “Abl.comp”), we observe a trend that more extensive searches (e.g., 150% or 200%) continue to offer gains. This confirms our intuition that more exhaustive data-mixture optimization can further boost downstream accuracy when sufficient compute is available.
@@ 6 Analysis
Effects of Proxy Model. Our method relies on a proxy model to rapidly score candidate mixtures. Intuitively, larger proxy models should better approximate the performance of the final (larger) target model. We test three proxy sizes: 62M, 132M, and 350M parameters. From Table 3 (rows under “Abl.proxy”), as we increase the proxy model from 62M to 350M, the average score improves from 60.11 to 60.41. Although the gains are not dramatic, they consistently favor using the largest feasible proxy model. This shows that a stronger proxy—closer in capacity to the target—achieves more accurate gradient estimates of mixture quality.
Effects of Number of Clusters. In our method, we employ a hierarchical clustering procedure. Specifically, we first group all data into $K_{init}$ clusters, perform a filtering step, and then regroup these clusters into $K_{enhanced}$ super-clusters. In this section, we explore the robustness of our data-mixture method and investigate its sensitivity to the number of clusters. Hence, we experiment with different values of $K_{init}$ (48, 64, 100, 1000, 2000) and $K_{enhanced}$ (15, 21, 30). The results in Table 3 (rows under “Abl.clus”) show that performance improves as $K_{init}$ increases from 48 to 100 and declines when $K_{init}$ increases from 1000 to 2000. Overall, our method is not particularly sensitive to the number of clusters, demonstrating the robustness of our approach. It is worth mentioning that if $K_{init}$ exceeds 2000 given the dataset size, the clustering becomes overly fine-grained and thus too dispersed. Likewise, if $K_{enhanced}$ is set too high, it requires more compute for sampling, increasing the overall cost of the data search process.
Effects of Initialization. We compare how different initialization schemes for the mixture weights affect performance. We experiment with a simple random initialization versus a Dirichlet-based initialization that biases weights to be more evenly spread at the start. Table 3 (rows under “Abl.init”) shows that Dirichlet initialization achieves a slightly higher average score (60.41%) than random (60.21%). The performances are comparable, suggesting the robustness of our data mixing approach, which is largely insensitive to the choice of initialization.
Evolution of Cluster Weights. The data mixture weights are important to understand the impact of different clusters, so we closely examine how they evolve across iterations. Figure 8 (a) presents the weights discovered by our search process for the 350M proxy model in the general reasoning domain. As shown, most clusters have minimal or no contribution (weights close to 0.00), while a few clusters play a significant role, with their weights changing across iterations. Among them, C18, C19, and C21 initially have high weights, but C19 and C21 exhibit a decreasing trend, suggesting their diminishing impact. Conversely, C8 and C9 become more relevant in later iterations, with their weights increasing in Iteration 3 (C8: 0.13, C9: 0.18), highlighting an adaptation in feature importance. More discussion about the evoluation of different cluster weights are shown in Appendix A.4.
Figure 6: Weight analysis of ClimbMix across iterations.
Analysis of Final Weights Furthermore, we analyzed the weights of the final data mixtures. From Figure 8 (a) , for the general reasoning task, C8, C9, C18, and C19 account for the majority of the weight. As shown in A.3, C8, C9, and C19 exhibit a high degree of correlation with general reasoning. Moreover, when analyzing the topics of these four clusters (4), we find that they collectively form a diverse distribution. In addition, we analyzed the importance of different clusters across domains on MMLU. As shown in Figures 8 (b), (c)), and (d), certain clusters play a crucial role in specific domains. For example, C7, C11, and C19 are particularly important for the humanities domain, while C7 and C8 are highly influential in the STEM domain. These findings highlight how different clusters contribute uniquely to various domains, providing deeper insights into domain-specific feature significance. We are also curious about the similarities and differences in the weights discovered by the large proxy model and the small proxy model. To explore this, we compared Figure 8 (a) and (e) and observed that they share similar important features, such as C8, C9, C18, and C19, although the assigned weights vary between the models. This insight suggests that we can leverage a smaller 62M proxy model for further experiments, reducing computational costs while retaining key structural patterns. The experimental results are presented in Appendix A.6. Notably, the weights appear sparse because, during the sampling process, we intentionally bias towards sparse weights. This approach effectively amplifies important clusters while filtering out less significant ones, enhancing the clarity of key features. In addition, we also investigate the relationship between clusters and downstream task performance in A.3.
@@
7 ClimbMix: New Pre-training Data
@@
Based on the insights obtained from our explorations above, we further apply CLIMB to two existing datasets: Nemotron-CC [8] and smollm-corpus [9], with the goal of constructing a powerful new pre-training dataset. Specifically, we first combine Nemotron-CC and smollm-corpus, and then employ our proposed CLIMB-clustering method to semantically reorganize and filter this combined dataset into 20 distinct clusters, leading to a 1.2-trillion-token high-quality corpus, named ClimbLab. Subsequently, we utilize CLIMB-search to identify an optimal data mixture from these clusters. Using this optimal mixture, we further extract a 400-billion-token high-quality dataset named ClimbMix. We train a 1B model from scratch with ClimbMix and evaluate its performance relative to models pretrained on other datasets under the same token budget. The results, illustrated in Figure 1, indicate that models trained on ClimbMix significantly outperform those trained on existing datasets. The optimal data mixture weights identified by CLIMB is shown in Figure 6. We note that in the previous continuous pre-training setting, a few domains accounted for the majority of the weight. However, since the experiments here are conducted under a pre-training-from-scratch setting, a more balanced cluster distribution is required compared to continuous pre-training. This difference arises because continuous pre-training provides a strong foundation, allowing the model to focus primarily on learning a few important domains, whereas pre-training from scratch necessitates more diverse data coverage. Finally, we publicly release these two datasets: the filtered 1.2-trillion-token dataset organized into 20 semantic clusters as a research playground for further data-mixture studies, and the optimized 400-billion-token ClimbMix dataset for efficient pre-training.
@@ 8 Conclusion
Appendix A Appendix
A.1 Experimental Settings
A.1.1 Baselines
@@
A.1.2 Data
For the source data, we employ Nemotron-CC [8] – a large dataset filtered from Common Crawl. It divides all the data into 20 buckets based on data quality annotation and we use the subset from the highest-quality bucket that consists of around 500 billion tokens. The hierarchical clustering of this subset results in 16 super-clusters (Cluster 1 to 16). To improve on data realism, we further incorporate five existing high-quality clusters. Three of these are sourced from the smollm-corpus [9]: Cosmopedia (Cluster 18), FineWeb-Edu (Cluster 19), and Python-Edu (Cluster 20). The remaining two clusters (Cluster 17 and Cluster 21) consist of additional academic data. In total, the source dataset consists of approximately 800 billion tokens distributed across 21 clusters.
@@ A.1.3 Model
A.2 Topics of Clusters
@@ 21 Government, Law, Scientific Revolution, Music, Literature
A.3 Relationship between Clusters and Downstream Tasks
@@
A.4 Evolution of Cluster Weights
@@
A.5 Analysis of Final Weights
Furthermore, we analyzed the weights of the final data mixtures. From Figure 8 (a) , for the general reasoning task, C8, C9, C18, and C19 account for the majority of the weight. As shown in A.3, C8, C9, and C19 exhibit a high degree of correlation with general reasoning. Moreover, when analyzing the topics of these four clusters (4), we find that they collectively form a diverse distribution.
In addition, we analyzed the importance of different clusters across domains on MMLU. As shown in Figures 8 (b), (c)), and (d), certain clusters play a crucial role in specific domains. For example, C7, C11, and C19 are particularly important for the humanities domain, while C7 and C8 are highly influential in the STEM domain. These findings highlight how different clusters contribute uniquely to various domains, providing deeper insights into domain-specific feature significance. We are also curious about the similarities and differences in the weights discovered by the large proxy model and the small proxy model. To explore this, we compared Figure 8 (a) and (e), and observed that they share similar important features, such as C8, C9, C18, and C19, although the assigned weights vary between the models. This insight suggests that we can leverage a smaller 62M proxy model for further experiments, reducing computational costs while retaining key structural patterns. The experimental results are presented in Appendix A.6. Notably, the weights appear sparse because, during the sampling process, we intentionally bias towards sparse weights. This approach effectively amplifies important clusters while filtering out less significant ones, enhancing the clarity of key features. In addition, we also investigate the relationship between clusters and downstream task performance in A.3.
A.6 Experiments with 62M proxy model
@@
A.7 Effects of Predictor

### v2

@@ 5 Analysis
Effects of Search Compute Budget. In the main experiments, we fix our total search budget (total compute) at 100%. Concretely, we perform three iterations of search with 64, 32, and 16 candidates evaluated in iterations 1, 2, and 3, respectively, giving a total of 112 searches. To understand how scaling search computes helps, we compare runs with greater total numbers of searches (e.g., 168, 224). Increasing the total number of searches allows the search procedure to more thoroughly explore possible data‐mixture candidates each iteration. As shown in Table 3 (rows under “Abl.comp”), we observe a trend that more extensive searches (e.g., 150% or 200%) continue to offer gains. This confirms our intuition that more exhaustive data‐mixture optimization can further boost downstream accuracy when sufficient compute is available.
@@
Analysis of Final Weights Furthermore, we analyzed the weights of the final data mixtures. From Figure 7 (a) , for the general reasoning task, C8, C9, C18, and C19 account for the majority of the weight. As shown in Appendix D.2, C8, C9, and C19 exhibit a high degree of correlation with general reasoning. Moreover, when analyzing the topics of these four clusters (Table 4), we find that they collectively form a diverse distribution. More detailed analysis is shown in Appendix D.8.
@@
In addition, we also discussed the topics of clusters, the relationship between clusters and downstream tasks, the effects of proxy models, the effects of the number of clusters, the effects of initialization, the effects of compute allocation, and the evolution of cluster weights in Appendix D.
6 Nemotron-ClimbLab and Nemotron-ClimbMix: New SOTA Pre-training Data
@@
Based on the insights obtained from our explorations above, we further apply CLIMB to two existing datasets: Nemotron-CC (su2024nemotron, ) and smollm-corpus (benallal2024smollmcorpus, ), with the goal of constructing a powerful new pre-training dataset. Specifically, we first combine Nemotron-CC and smollm-corpus, and then employ our proposed CLIMB-clustering method to semantically reorganize and filter this combined dataset into 20 distinct clusters, leading to a 1.2-trillion-token high-quality corpus, named Nemotron-ClimbLab. Subsequently, we utilize CLIMB-search to identify an optimal data mixture from these clusters. Using this optimal mixture, we further extract a 400-billion-token high-quality dataset named Nemotron-ClimbMix. We train a 1B model from scratch with Nemotron-ClimbMix and evaluate its performance relative to models pretrained on other datasets under the same token budget. The results, illustrated in Figure 1, indicate that models trained on Nemotron-ClimbMix significantly outperform those trained on existing datasets, including Nemotron-CC (su2024nemotron, ), SmolLM (benallal2024smollmcorpus, ), DCLM-baseline (li2024datacomp, ), and FineWeb-Edu (penedo2024the, ). The optimal data mixture weights identified by CLIMB is shown in Figure 8. We note that in the previous continuous pre-training setting, a few domains accounted for the majority of the weight. However, since the experiments here are conducted under a pre-training-from-scratch setting, a more balanced cluster distribution is required compared to continuous pre-training. This difference arises because continuous pre-training provides a strong foundation, allowing the model to focus primarily on learning a few important domains, whereas pre-training from scratch necessitates more diverse data coverage. Finally, we publicly release these two datasets: the filtered 1.2-trillion-token dataset organized into 20 semantic clusters as a research playground for further data-mixture studies, and the optimized 400-billion-token Nemotron-ClimbMix dataset for efficient pre-training.
7 Related Work
Data Mixture for LLM Pre-training. The composition of pre-training datasets are critical in determining the generalization abilities of language models  (radford2019language, ; brown2020language, ; touvron2023llama, ). Typically, data mixtures like those in the Pile (gao2020pile, ), GLaM (du2022glam, ), and ROOTS (laurenccon2022bigscience, ) are crafted using manually defined rules, yet these heuristics lack standardization and transferability across different settings. SlimPajama-DC (shen2023slimpajama, ) systematically evaluated the influence of various predefined data configurations, yielding valuable insights. More recently, learning-based approaches such as DoReMi (xie2024doremi, ) and DoGE (fan2023doge, ) have introduced optimization techniques for domain proportions by iteratively refining training with reference and proxy models. While (chen2024skill, ) investigated data sequencing strategies through the lens of curriculum learning, our work focuses on the simultaneous integration of diverse data domains, emphasizing a distinct aspect of pre-training. The aforementioned methods show promise, but they require the dataset to already possess clear and natural domain distinctions. By contrast, we propose a novel approach that can automatically identify approximate domains from large amounts of web data and then find the optimal data mixture automatically. In parallel work, WebOrganizer (wettig2025organize, ) proposes using classifiers to annotate web-scale data with topic and format labels. In contrast, our clustering-based approach is more straightforward, and readily scalable, and we introduce an iterative optimization method to refine the data mixture.
Data Selection for Specific Domains. Beyond optimizing the overall pre-training data mixture (zhang2024harnessing, ; yu2024mates, ; wettigqurating, ; brandfonbrener2024color, ), selecting high-quality domain-specific data (ruder2017learning, ; xie2024efficient, ) is essential for improving model specialization during pre-training. Existing methods approach this challenge differently. DSIR (xie2023data, ) estimates relevance using hashed n-grams and resamples data to better match target domain distributions. CRISP (grangier2025taskadaptive, ) clusters the generalist dataset and samples these clusters according to their frequencies in the smaller specialist dataset. (kang2024get, ) propose to select data that nudges the pre-training distribution closer to the target distribution. Training dynamics-based selection leverages model learning behavior to guide data filtering, including S2L (yang2024smalltolarge, ), which clusters data based on loss trajectories to prioritize domain-relevant examples, and LESS (xia2024less, ), which selects instruction tuning data with the highest gradient similarity to a target task. Embedding-based filtering removes redundant (abbas2023semdedup, ) or low-quality data, with SCIP (yang2023decoding, ) applying synthetic corruptions for filtering and heuristic pruning (singh2024brevity, ) reducing noise from overrepresented long-text clusters. (shum2025predictive, ) proposes to select the data on which model losses are predictive of downstream abilities. While these approaches improve data quality for specialized domains, they often rely on predefined domain labels or heuristics, limiting their flexibility for large-scale pre-training. In contrast, our proposed framework, CLIMB, iteratively refines domain-relevant data mixtures without requiring explicit domain labels, making it more applicable to real-world pre-training data and easier to scale.
@@ 8 Conclusion
Appendix A Limitations
While our proposed CLIMB framework demonstrates strong performance and provides valuable insights into data mixture optimization, several limitations warrant further exploration.
First, although we mitigate computational overhead by leveraging lightweight proxy models during the iterative search process, training these proxy models still incurs non-negligible costs. We show that using lightweight models (e.g., 350M and 62M) is feasible, but further reducing the computational burden—perhaps through parameter-efficient tuning, distillation, or zero-shot evaluation strategies—remains an important direction for future work.
Second, our evaluation of domain-specific benefits is based on MMLU’s coarse-grained domain categories (e.g., STEM, Social Sciences), which may not fully reflect real-world applications. While our results highlight the potential of CLIMB for domain adaptation, we have not yet evaluated its effectiveness in high-stakes domains such as finance or healthcare, where data characteristics and requirements can differ substantially. We leave such real-world validation to future research.
Appendix B Societal Impacts
This work contributes to advancing the field of machine learning by proposing an automated framework for optimizing data mixtures in language model pre-training. Our method enables more efficient and scalable objective-aware pre-training by leveraging clustering and iterative search, reducing reliance on manually curated datasets.
From an ethical standpoint, our approach does not introduce new risks beyond those commonly associated with LLM training, such as biases in training data and potential misuse. However, optimizing the data mixture may lead to an overrepresentation of certain domains while underrepresenting others, potentially reinforcing knowledge disparities. While our method allows for more controlled and targeted data composition, future work should explore safeguards to ensure fairness and mitigate unintended biases in domain representation. Additionally, our method automates domain-aware data selection without requiring explicit human annotation, potentially reducing reliance on curated datasets. However, this also raises concerns about unintended biases in automatically clustered data. Ensuring transparency and interpretability in data selection remains a critical area for future exploration.
Appendix C Experimental Settings
C.1 Baselines
@@
C.2 Data
For the source data, we employ Nemotron-CC [8] – a large dataset filtered from Common Crawl. It divides all the data into 20 buckets based on data quality annotation, and we use the subset from the highest-quality bucket. The hierarchical clustering of this subset results in approximately 800 billion tokens distributed across 21 clusters.
@@ C.3 Model
C.4 Training Settings
For pre-training, we use AdamW optimizer and set the learning rate to 5e-5 for the stable stage and anneal it to 1e-5. We use a batch size of 2M tokens throughout the training process, utilizing 256 NVIDIA H100 GPUs. The training time of a single lightweight proxy model is approximately 45 GPU hours, while the training time of the large target model is around 6,400 GPU hours.
Appendix D Analysis
D.1 Topics of Clusters
@@ 21 Government, Law, Scientific Revolution, Music, Literature
D.2 Relationship between Clusters and Downstream Tasks
@@
D.3 Effects of Proxy Model.
Our method relies on a proxy model to rapidly score candidate mixtures. Intuitively, larger proxy models should better approximate the performance of the final (larger) target model. We test three proxy sizes: 62M, 132M, and 350M parameters. From Table 3 (rows under “Abl.proxy”), as we increase the proxy model from 62M to 350M, the average score improves from 60.11 to 60.41. Although the gains are not dramatic, they consistently favor using the largest feasible proxy model. This shows that a stronger proxy—closer in capacity to the target—achieves more accurate gradient estimates of mixture quality.
D.4 Effects of Number of Clusters.
In our method, we employ a hierarchical clustering procedure. Specifically, we first group all data into $K_{init}$ clusters, perform a filtering step, and then regroup these clusters into $K_{enhanced}$ super-clusters. In this section, we explore the robustness of our data‐mixture method and investigate its sensitivity to the number of clusters. Hence, we experiment with different values of $K_{init}$ (48, 64, 100, 1000, 2000) and $K_{enhanced}$ (15, 21, 30). The results in Table 3 (rows under “Abl.clus”) show that performance improves as $K_{init}$ increases from 48 to 100 and declines when $K_{init}$ increases from 1000 to 2000. Overall, our method is not particularly sensitive to the number of clusters, demonstrating the robustness of our approach. It is worth mentioning that if $K_{init}$ exceeds 2000 given the dataset size, the clustering becomes overly fine‐grained and thus too dispersed. Likewise, if $K_{enhanced}$ is set too high, it requires more compute for sampling, increasing the overall cost of the data search process.
D.5 Effects of Initialization.
We compare how different initialization schemes for the mixture weights affect performance. We experiment with a simple random initialization versus a Dirichlet‐based initialization that biases weights to be more evenly spread at the start. Table 3 (rows under “Abl.init”) shows that Dirichlet initialization achieves a slightly higher average score (60.41%) than random (60.21%). The performances are comparable, suggesting the robustness of our data mixing approach, which is largely insensitive to the choice of initialization.
D.6 Effects of Compute Allocation.
By default, we allocate our 100% total compute across three iterations in a 4:2:1 ratio (64:32:16). In principle, however, one could allocate compute to create either a “tall” search tree (more iterations but fewer searches per iteration) or a “fat” one (fewer iterations but more searches per iteration). Table 3 (rows under “Abl.allo”) compares several such allocations: 6:1, 4:2:1, and 2:2:1:1. We find that 4:2:1 yields the best overall average performance (60.41%). Having too few iterations (e.g., 6:1) can lead to suboptimal exploration in earlier iterations, while splitting too many iterations (2:2:1:1) spreads compute too thin across each iteration. Thus, balancing depth (number of iterations) and breadth (searches per iteration) proves key to robustly finding a good mixture.
D.7 Evolution of Cluster Weights
@@
D.8 Analysis of Final Weights
Furthermore, we analyzed the weights of the final data mixtures. From Figure 7 (a) , for the general reasoning task, C8, C9, C18, and C19 account for the majority of the weight. As shown in D.2, C8, C9, and C19 exhibit a high degree of correlation with general reasoning. Moreover, when analyzing the topics of these four clusters (4), we find that they collectively form a diverse distribution.
In addition, we analyzed the importance of different clusters across domains on MMLU. As shown in Figures 7 (b), (c), and (d), certain clusters play a crucial role in specific domains. For example, C7, C11, and C19 are particularly important for the humanities domain, while C7 and C8 are highly influential in the STEM domain. These findings highlight how different clusters contribute uniquely to various domains, providing deeper insights into domain-specific feature significance. We are also curious about the similarities and differences in the weights discovered by the large proxy model and the small proxy model. To explore this, we compared Figure 7 (a) and (e), and observed that they share similar important features, such as C8, C9, C18, and C19, although the assigned weights vary between the models. This insight suggests that we can leverage a smaller 62M proxy model for further experiments, reducing computational costs while retaining key structural patterns. The experimental results are presented in Appendix D.9. Notably, the weights appear sparse because, during the sampling process, we intentionally bias towards sparse weights. This approach effectively amplifies important clusters while filtering out less significant ones, enhancing the clarity of key features. In addition, we also investigate the relationship between clusters and downstream task performance in D.2.
D.9 Experiments with 62M proxy model
@@
Figure 8: Weight analysis of Nemotron-ClimbMix across iterations.
D.10 Effects of Predictor